With a fact that at STP a mole of a gas occupies 22.4 L, calculate the entropy change
of mixing of 2.8 litres of O2 with 19.6 litres of H2 at STP. Assume the gases to behave
ideally
1
Expert's answer
2020-04-23T13:08:04-0400

The Mixing of Ideal Gases:


is the mole fraction!


n (O2) = 2.8 / 22.4 = 0.125 mol

n (H2) = 19.6 / 22.4 = 0.875 mol

Total: 0.125 mol + 0.875 mol = 1 mol

mole fraction = 0.125 (O2)

mole fraction = 0.875 (H2)


Entropy change = - 1 * 8.314 * (0.125 ln(0.125) + 0.875 ln (0.875))

Entropy change = +3.13 J/K